# Safe and Optimal N-Spacecraft Swarm Reconfiguration in Non-Keplerian Cislunar Orbits

### v1

@@ 2.2 Absolute Dynamics of RMBPs
By convention, the equations of motion in RMBPs are non-dimensionalized by the (a) Earth-Moon distance (length), (b) the sum of the primary mass (mass), and (c) time, so that the gravitational constant of the Earth-Moon system becomes unity. For all these dynamical models, the equations of motion in the barycenter synodic frame are written as follows:
@@
where $\Upsilon$ is a pseudo-potential function; $\boldsymbol{\omega}_{b/i}=[\omega^{x}_{b/i},\omega^{y}_{b/i},\omega^{z}_{b/i}]$ is the angular velocity of the barycenter rotating frame with respect to the inertial frame; and $\boldsymbol{r}_{b_{1}c}=[x,y,z]$ is the position of the chief with respect to $B_{1}$. For example, the pseudo-potential of the CR3BP is given by
@@
where $\mu=\dfrac{m_{m}}{m_{e}+m_{m}}$ is the mass ratio of the two primaries. The CR3BP assumption leads to $\boldsymbol{\omega}_{b/i}=[0,0,1]$ (i.e., $[\dot{\boldsymbol{\omega}}_{b/i}]_{\mathcal{B}}=\boldsymbol{0}$). The pseudo-potenitals of ER3BP and BCR4BP are derived in a similar manner and are shown in Appendix A. For the ER3BP, the angular velocity and the Earth-Moon distance are the time-varying parameters. Furtheremore, the BCR4BP assumes the Sun’s position in the barycenter synodic frame to be $\boldsymbol{r}_{b_{1}s}=[x_{s},y_{x},z_{s}]=[r_{b_{1}s}\cos\theta_{s},r_{b_{1}s}\sin\theta_{s},0]$, which rotates around the barycenter at a constant angular velocity $\boldsymbol{\omega}_{s}$ with its non-dimensionalized mass $\mu_{s}=\dfrac{m_{s}}{m_{e}+m_{m}}$, as shown in Fig. 1(b).
2.3 Relative Dynamics in the TNW frame
Following the derivations of the relative dynamics in the CR3BP, ER3BP, and BCR4BP within the LVLH frame [16, 18], this subsection develops nonlinear and linear relative dynamics in R3BP in the TNW frame. While this derivation assumes that the spacecraft swarm is anchored to the Moon, the extension to orbits around other anchoring points (e.g., libration points) is readily available [17].
A key difference in the derivation in the TNW frame to that in the LVLH frame [16] is the analytical expression of the angular velocity of the TNW frame with respect to the Moon synodic frame $\boldsymbol{\omega}_{v/m}$ and angular acceleration $[\dot{\boldsymbol{\omega}}_{v/m}]_{\mathcal{V}}$, which are derived as
@@ 2.3 Relative Dynamics in the TNW frame
$\dot{\omega}^{x}_{v/m}$ $=-\dfrac{\dot{h}}{h^{2}}((\boldsymbol{r}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot\hat{\boldsymbol{k}})+\dfrac{1}{h}\left(([\dot{\boldsymbol{r}}]_{\mathcal{M}}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot\hat{\boldsymbol{k}}+(\boldsymbol{r}\times[\dddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot\hat{\boldsymbol{k}}+(\boldsymbol{r}\times[\ddot{\boldsymbol{r}}]_{\mathcal{M}})\cdot[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}}\right)$ (5b)
$\dot{\omega}^{y}_{v/m}$ $=\dfrac{\dot{v}}{v^{2}}([\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}})-\frac{1}{v}([\dddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot\hat{\boldsymbol{k}}+[\ddot{\boldsymbol{r}}]_{\mathcal{M}}\cdot[\dot{\hat{\boldsymbol{k}}}]_{\mathcal{M}})$ (5c)
@@
$\dot{\boldsymbol{x}}=\boldsymbol{A}(\boldsymbol{r}(t))\boldsymbol{x}=\begin{bmatrix}\boldsymbol{0}_{3}&\boldsymbol{I}_{3}\\ \boldsymbol{A}_{\dot{\boldsymbol{\rho}}\boldsymbol{\rho}}&-2[\boldsymbol{\omega}_{v/i}]^{\times}\end{bmatrix}\boldsymbol{x},\quad\boldsymbol{x}=\begin{bmatrix}\boldsymbol{\rho}\\ [\dot{\boldsymbol{\rho}}]_{\mathcal{V}}\end{bmatrix}$ (7a)
where $\boldsymbol{A}_{\dot{\boldsymbol{\rho}}\boldsymbol{\rho}}=-[\dot{\boldsymbol{\omega}}_{v/i}]_{\mathcal{V}}^{\times}-\left([\boldsymbol{\omega}_{v/i}]_{\mathcal{V}}^{\times}\right)^{2}-\mu\frac{\partial}{\partial\boldsymbol{r}}\left[\frac{\boldsymbol{r}}{r^{3}}\right]-(1-\mu)\frac{\partial}{\partial\boldsymbol{r}}\left[\frac{\boldsymbol{r}+\boldsymbol{r}_{em}}{\left\|\boldsymbol{r}+\boldsymbol{r}_{em}\right\|^{3}}\right]-\mu_{s}\frac{\partial}{\partial\boldsymbol{r}}\left[\frac{\boldsymbol{r}+\boldsymbol{r}_{sm}}{\left\|\boldsymbol{r}+\boldsymbol{r}_{sm}\right\|^{3}}\right].$ (7b)
Note that ${\boldsymbol{\omega}}_{v/i}={\boldsymbol{\omega}}_{v/m}+{\boldsymbol{\omega}}_{m/i}$ and $[\dot{\boldsymbol{\omega}}_{v/i}]_{\mathcal{V}}=[\dot{\boldsymbol{\omega}}_{v/m}]_{\mathcal{V}}+[\dot{\boldsymbol{\omega}}_{m/i}]_{\mathcal{M}}-\boldsymbol{\omega}_{v/m}\times\boldsymbol{\omega}_{m/i}$ [16].
@@ 3 Bounded Relative Motion in Multi-body Problems
In this section, a thorough comparison is presented between natural bounded relative motions in the RP2BP and the RMBPs. Due to dynamics’ nonlinear and chaotic nature, the computation of bounded relative motions in the RMBPs has heavily relied on numerical continuation methods [33, 34]. However, by analyzing the linearized dynamical system and the fundamental solution matrix [40] offers a geometric insight into the relative motion in RMBPs. Available dynamical modes in two dynamical systems are first presented, followed by the introduction of the LTC [26] that leverages the dynamical mode that excites the bounded relative motion in the RMBPs.
@@ 3.1 Eigensystem in the Vicinity of Periodic Orbits
[…]
where a plant matrix $A(\tau)=\dfrac{\partial f}{\partial\tau}$ is the linearized dynamics of a system evaluated at the reference orbit. The eigensystem of the monodromy matrix defines the linear stability of the states near the periodic solution.
Also, consider a linear mapping from the Cartesian state perturbation $[\boldsymbol{\rho}^{\top},\dot{\boldsymbol{\rho}}^{\top}]^{\top}$ from a reference periodic solution to a set of new state representation $\boldsymbol{\kappa}(t)\in\mathbb{R}^{6}$ as follows:
$\boldsymbol{\kappa}(t)=\Psi^{-1}(t)\begin{bmatrix}\boldsymbol{\rho}(t)\\ \dot{\boldsymbol{\rho}}(t)\end{bmatrix}.$ (9)
@@
Note that $\Psi(t_{0}+T)=\Psi(t_{0})\forall t_{0}$ because of the periodicity of the reference solution.
@@ 3.1.1 Eigensystem in the R2BP
In Keplerian dynamics, it is known that all six eigenvalues of the monodromy matrix are always one. The dynamical mode associated with this eigenvalue is called a central mode. The expression of the monodromy matrix for circular orbits (i.e., HCW equation [3]) is analytically available (cf. Appendix D). An analytical expression is not available for relative motion in eccentric orbits due to the time-varying dynamics [4], although the same eigensystem can be obtained. The five unique eigenvectors are unit vectors of all six directions except for a mode in the along-track direction, as shown in Fig. 2(a). This eigensystem induces a periodic relative motion from any initial condition, except for the along-track direction. This component exhibits drift in the case of differences in specific mechanical energy (or semi-major axis) as approximated by the linearization [5]. Therefore, based on Eq. 10, $M_{\boldsymbol{\kappa}}$ retains six unit eigenvalues (i.e., periodicity of $\kappa(t)$) for an arbitrary linear invertible map $\Psi(t)$. Thus, a set of invariant integration constants $\kappa(t)=\kappa$ can be defined with $\Psi(t)$ being a fundamental matrix solution to the variational equation, which leads to the well-known ROE theory [5]. Note that this eigensystem provides five degrees of freedom in the design space of bounded relative orbits.
@@
When additional perturbations (i.e., PR2BP) are considered, some eigenvalues of the monodromy matrix perturb from unity. Due to the additional conditions required for periodicity, the solution set for periodic relative motion is reduced. Previous works address bounded relative orbit designs that cancel a secular drift caused by the $J_{2}$-perturbation [5, 36].
@@ 3.1.2 Eigensystem in the RMBP
Similar to an elliptic orbit, the local linearization of the dynamics of RMBPs contains time-varying components, preventing a closed-form expression of the monodromy matrix. Instead, the monodromy matrix can be numerically integrated (cf. Eq. (8)) concurrently with the periodic solution (cf. Eqs. (36a), (38a), and (39a)) for the time interval $[t_{0},t_{0}+T]$, using the plant matrix $\boldsymbol{A}(t)$. This system becomes a 42-dimensional ordinary differential equation.
Generally, not all eigenvalues of the monodromy matrix computed along a periodic solution of RMBP are one. Therefore, the entire set of bounded relative periodic solutions around the absolute periodic solution is severely constrained. One trivial solution is a set of relative motions between spacecraft on the same periodic absolute orbit but with phase offsets. Besides, an absolute periodic orbit in the RMBPs may admit up to two oscillatory modes in the corresponding monodromy matrix, characterized by a conjugate pair of complex unimodular eigenvalues [14, 26]. When the oscillatory mode is excited, an initial relative state within the corresponding eigenspace evolves into a quasi-periodic motion; after one period $T$, the relative position and velocity do not repeat but remain confined within a rotating ellipse, known as an invariant circle. This structure is illustrated in Fig.  2(b) as dotted ellipses on the eigenspace. The rotational displacement in a period is called rotation number [33, 34], determined as $\varrho:=\tan^{-1}(\lambda_{r}/\lambda_{i})$, where $\lambda_{r}+\lambda_{i}i$ is the eigenvalue corresponding to the oscillatory mode.
@@
Figure 3: Periodic and quasi-periodic absolute motion in the R2BP and RMBP.
When resolved in an absolute frame, a QPRO constructed based on the oscillatory mode in the linearized dynamics is equivalent to an absolute QPO that envelopes a reference orbit under the original nonlinear dynamics. The QPO goes along the surface of a two-dimensional Quasi-Periodic Invariant Torus (QPIT), which is a stable geometric structure along which solutions evolve without altering their overall shape, as shown in Fig. 3(b) This bounded motion is ensured because each point on the torus surface follows the vector field of the system’s equations of motion. Any position and velocity ($\mathbb{R}^{6}$) on the surface of a QPIT can be characterized by a two-dimensional phase vector $\boldsymbol{\theta}=[\theta_{0},\theta_{1}]\in\mathbb{T}^{2}$. Consequently, a motion initiated from a point on the torus can be viewed as an evolution of $\boldsymbol{\theta}$. The progression along a periodic absolute orbit is characterized by $\theta_{0}$. In contrast, $\theta_{1}$ represents the phase within an invariant curve, defined as a closed curve where quasi-periodic solutions are confined under the $T$-stroboscopic map. As shown in Fig. 3(a), a periodic absolute orbit is essentially a one-dimensional torus, parameterized by $\boldsymbol{\theta}=\theta_{0}\in\mathbb{T}$. The QPIT resolved in the co-moving frame is termed Quasi-Periodic Relative Invariant Torus (QPRIT). Thus, an invariant circle is equivalent to the first-order approximation of the invariant curve (cf. Fig. 3(b)) in a co-moving frame [33], leading to two degrees of freedom in the relative orbit design in each oscillatory mode. Furthermore, the rotation number $\varrho$ is the first-order approximation of the drift in $\theta_{1}$ on the invariant curve.
@@ 3.2 Local Toroidal Coordinates
Local Toroidal Coordinates (LTC) [26] are coordinate systems on a time-varying frame that project the dynamics onto the oscillatory eigenspace of the RMBPs. Instead of a full linear mapping to integration constants, a mode-isolating transformation is employed to emphasize the dynamical mode that provides a (meaningful) quasi-periodic relative motion. It is assumed in this paper that the chief is on a periodic orbit of RMBPs that contains at least one oscillatory mode.
@@
$\boldsymbol{\psi}(t)=\varepsilon\left(\boldsymbol{w}_{r}(t)\cos\theta+\boldsymbol{w}_{i}(t)\sin\theta\right),\quad\boldsymbol{w}_{r}=\begin{bmatrix}\boldsymbol{r}_{r}(t)\\ \boldsymbol{v}_{r}(t)\end{bmatrix},\boldsymbol{w}_{i}=\begin{bmatrix}\boldsymbol{r}_{i}(t)\\ \boldsymbol{v}_{i}(t)\end{bmatrix},$ (11)
Figure 4: Nonsingular local toroidal coordinates $[\alpha,\beta,h]$ and geometric local toroidal coordinates $[\varepsilon,\theta,h]$
@@
With this in mind, the local toroidal frame is defined based on a time-varying non-orthogonal and non-unimodular basis triad $\{\boldsymbol{r}_{r}(t),\boldsymbol{r}_{i}(t),\hat{\boldsymbol{n}}(t)\}$, where $\hat{\boldsymbol{n}}(t)$ is a unimodular vector parallel to $\boldsymbol{r}_{r}(t)\times\boldsymbol{r}_{i}(t)$. In order to achieve a unique and consistent representation of an eigenvector and to construct a triad from a local eigensystem, basis vectors at the fixed point at the initial epoch $\{\boldsymbol{r}_{r}(t_{0}),\boldsymbol{r}_{i}(t_{0})\}$ are normalized and rotated via singular value decomposition (SVD) so that these vectors are orthogonal and $\boldsymbol{r}_{r}(t_{0})$ becomes unimodular. For details of the normalization process, see Ref. 22. Once normalized at the fixed point, the evolution of $\boldsymbol{w}$ can be propagated by the linear dynamics
$\dot{\boldsymbol{w}}(t)=\boldsymbol{A}(t)\boldsymbol{w}(t),$ (12)
where the plant matrix is obtained from Eq. (7). Therefore, the basis vectors $\{\boldsymbol{r}_{r}(t),\boldsymbol{r}_{i}(t),\hat{\boldsymbol{n}}(t)\}$ are generally neither orthogonal (nor unimodular for $\{\boldsymbol{r}_{r},\boldsymbol{r}_{i}\}$) except for the initial triad at the fixed point.
Within the local toroidal frame, the relative position of the deputy to the chief is written as $\boldsymbol{\rho}=\alpha\boldsymbol{r}_{r}+\beta\boldsymbol{r}_{i}+h\hat{\boldsymbol{n}}$. Based on this, nonsingular LTC is defined as $\boldsymbol{\zeta}_{\text{ns}}:=[\alpha,\beta,h,\dot{\alpha},\dot{\beta},\dot{h}]^{\top}$, where time derivatives are taken in the local toroidal frame. using Eq. (9), the mapping from the nonsingular LTC $\boldsymbol{\zeta}_{\text{ns}}$ to the relative position and velocity expressed in the Cartesian coordinates $[\boldsymbol{\rho}^{\top},\dot{\boldsymbol{\rho}}^{\top}]^{\top}$ is related as
$\boldsymbol{\zeta}_{\text{ns}}=T(t)^{-1}\begin{bmatrix}\boldsymbol{\rho}\\ \dot{\boldsymbol{\rho}}\end{bmatrix},\quad T(t):=\begin{bmatrix}R(t)&\boldsymbol{0}_{3}\\ R^{\prime}(t)&R(t)\end{bmatrix}\Leftrightarrow T^{-1}(t)=\begin{bmatrix}R(t)^{-1}&\boldsymbol{0}_{3}\\ -R(t)^{-1}R^{\prime}(t)R(t)^{-1}&R(t)^{-1}\end{bmatrix},$ (13)
where $R(t):=[\boldsymbol{r}_{r}(t),\boldsymbol{r}_{i}(t),\hat{\boldsymbol{n}}(t)]$ and $R^{\prime}(t):=[\boldsymbol{v}_{r}(t),\boldsymbol{v}_{i}(t),\dot{\hat{\boldsymbol{n}}}(t)]$. The time derivative of the unit normal vector $\hat{\boldsymbol{n}}(t)$ is computed as
$\dot{\hat{\boldsymbol{n}}}(t)=(\boldsymbol{v}_{r}(t)\times\boldsymbol{r}_{i}(t)+\boldsymbol{r}_{r}(t)\times\boldsymbol{v}_{i}(t))-(\boldsymbol{n}(t)^{\top}(\boldsymbol{v}_{r}(t)\times\boldsymbol{r}_{i}(t)+\boldsymbol{r}_{r}(t)\times\boldsymbol{v}_{i}(t)))\boldsymbol{n}(t).$ (14)

### v2

@@ 2.3 Dynamics of RMBPs
By convention, the equations of motion in the Earth-Moon system are non-dimensionalized by the (a) Earth-Moon distance (length), (b) the sum of the primary mass (mass), and (c) angular rate of primaries around the Earth-Moon barycenter (time), so that the gravitational constant of the Earth-Moon system becomes unity. For all these dynamical models, the equations of motion resolved in the barycenter synodic frame are written as follows:
@@
where $\Upsilon$ is a pseudo-potential function; $\boldsymbol{\omega}_{b/i}$ is the angular velocity of the barycenter synodic frame with respect to the inertial frame, whose components in the barycenter synodic frame are given as $\omega^{x}_{b/i},\omega^{y}_{b/i},\omega^{z}_{b/i}]$; and $\boldsymbol{r}_{b_{1}c}$ is the position of the chief with respect to $B_{1}$, whose components in the barycenter synodic frame are given as $[x,y,z]$. For example, the pseudo-potential of the CR3BP is given by
@@
where $\mu=\dfrac{m_{m}}{m_{e}+m_{m}}$ is the mass ratio of the two primaries. The CR3BP assumption leads to $\omega^{x}_{b/i},\omega^{y}_{b/i},\omega^{z}_{b/i}]=[0,0,1]$ (i.e., $\dot{\boldsymbol{\omega}}_{b/i}]_{\mathcal{B}}=\boldsymbol{0}$). The pseudo-potenitals of ER3BP and BCR4BP are derived in a similar manner and are shown in Appendix A. For the ER3BP, the angular velocity and the Earth-Moon distance are the time-varying parameters. The BCR4BP parameterizes the Sun’s position in the barycenter synodic frame $\boldsymbol{r}_{b_{1}s}$ as a function of $\theta_{s}$, which is the angle between the Earth-Moon line and the Sun-$B_{1}$ line, as shown in Fig. 1(b). Within the barycenter synodic frame, the components of $\boldsymbol{r}_{b_{1}s}$ are described as $x_{s},y_{x},z_{s}]=[r_{b_{1}s}\cos\theta_{s},r_{b_{1}s}\sin\theta_{s},0]$, which rotates around the Earth-Moon barycenter at a constant angular velocity $\boldsymbol{\omega}_{s}$ with its non-dimensionalized mass $\mu_{s}=\dfrac{m_{s}}{m_{e}+m_{m}}$.
2.4 Relative Dynamics in the VNB
frame
Following the derivations of the relative dynamics in the CR3BP, ER3BP, and BCR4BP within the LVLH frame [19, 21], this subsection develops nonlinear and linear relative dynamics in R3BP in the VNB frame. While this derivation assumes that the spacecraft swarm is anchored to the Moon, the extension to orbits around other anchoring points (e.g., libration points) is readily available [20].
A key difference in the derivation in the VNB frame to that in the LVLH frame [19] is the analytical expression of the angular velocity of the VNB frame with respect to the Moon synodic frame $\boldsymbol{\omega}_{v/m}$ and angular acceleration $\dot{\boldsymbol{\omega}}_{v/m}]_{\mathcal{V}}$, which are derived as
@@ 2.4 Relative Dynamics in the VNB
[…]
@@
$\dot{\boldsymbol{x}}=\boldsymbol{A}(\boldsymbol{r}(t))\boldsymbol{x}=\begin{bmatrix}\boldsymbol{0}_{3}&\boldsymbol{I}_{3}\\ \boldsymbol{A}_{\dot{\boldsymbol{\rho}}\boldsymbol{\rho}}&-2[\boldsymbol{\omega}_{v/i}]^{\times}\end{bmatrix}\boldsymbol{x},\quad\boldsymbol{x}=\begin{bmatrix}\boldsymbol{\rho}\\ [\dot{\boldsymbol{\rho}}]_{\mathcal{V}}\end{bmatrix}$, (8a)
$\boldsymbol{A}_{\dot{\boldsymbol{\rho}}\boldsymbol{\rho}}=-[\dot{\boldsymbol{\omega}}_{v/i}]_{\mathcal{V}}^{\times}-\left([\boldsymbol{\omega}_{v/i}]_{\mathcal{V}}^{\times}\right)^{2}-\mu\frac{\partial}{\partial\boldsymbol{r}}\left[\frac{\boldsymbol{r}}{r^{3}}\right]-(1-\mu)\frac{\partial}{\partial\boldsymbol{r}}\left[\frac{\boldsymbol{r}+\boldsymbol{r}_{em}}{\left\|\boldsymbol{r}+\boldsymbol{r}_{em}\right\|^{3}}\right]-\mu_{s}\frac{\partial}{\partial\boldsymbol{r}}\left[\frac{\boldsymbol{r}+\boldsymbol{r}_{sm}}{\left\|\boldsymbol{r}+\boldsymbol{r}_{sm}\right\|^{3}}\right].$ (8b)
Note that ${\boldsymbol{\omega}}_{v/i}={\boldsymbol{\omega}}_{v/m}+{\boldsymbol{\omega}}_{m/i}$ and $\dot{\boldsymbol{\omega}}_{v/i}]_{\mathcal{V}}=[\dot{\boldsymbol{\omega}}_{v/m}]_{\mathcal{V}}+[\dot{\boldsymbol{\omega}}_{m/i}]_{\mathcal{M}}-\boldsymbol{\omega}_{v/m}\times\boldsymbol{\omega}_{m/i} [$19].
@@ 3 Bounded Relative Motion in Multi-body Problems
In this section, a thorough comparison is presented between natural bounded relative motions in the RP2BP and the RMBPs. Due to nonlinear and chaotic nature of dynamics, the computation of bounded relative motions in the RMBPs has heavily relied on numerical continuation methods [28, 29]. However, by analyzing the linearized dynamical system and the fundamental solution matrix [35] offers a geometric insight into the relative motion in RMBPs. Available dynamical modes in two dynamical systems are first presented, followed by the introduction of the LTC [40] that leverages the dynamical mode that excites the bounded relative motion in the RMBPs. To analyze the dynamics, this section focuses on the derivation in the VNB frame as a specific choice of the local frame. However, this can be readily replaced with the LVLH frame or other local frames.
@@ 3.1 Eigensystem in the Vicinity of Periodic Orbits
$M(t_{0}):=\Phi(t_{0},t_{0}+T)=\Phi(t_{0},t_{0})+\int_{0}^{T}A(\tau)\Phi(t_{0},t_{0}+\tau)d\tau,\quad\Phi(t_{0},t_{0})=\boldsymbol{I}_{6},$ (9)
where a plant matrix $A(\tau)=\dfrac{\partial f}{\partial\tau}$ is the linearized dynamics of a system evaluated at the reference orbit. The eigensystem of the monodromy matrix defines the linear stability of the perturbed states near the reference orbit.
Also, consider a linear mapping from the Cartesian state perturbation $\boldsymbol{x}=[\boldsymbol{\rho}^{\top},[\dot{\boldsymbol{\rho}}]_{\mathcal{V}}^{\top}]^{\top}$ from the reference orbit to a set of new state representation $\boldsymbol{\kappa}(t)\in\mathbb{R}^{6}$ as follows:
$\boldsymbol{\kappa}(t)=\Psi^{-1}(t)\boldsymbol{x}.$ (10)
@@
Note that $\Psi(t_{0}+T)=\Psi(t_{0})\forall t_{0}$ because of the periodicity of the reference orbit.
@@ 3.1.1 Eigensystem in the R2BP
Periodic orbits in the Keplerian dynamics are integrable and possess six first integrals, which implies that all six eigenvalues of the monodromy matrix are one, with five linearly independent eigenvectors [22, 6]. The dynamical mode associated with the unity eigenvalue is called a central mode. The five eigenvectors are unit vectors of all six directions except for a mode in the along-track direction, as shown in Fig. 2(a). This eigensystem induces periodic relative motion from any initial condition, except for the along-track direction, where a difference in specific mechanical energy induces drift [5]. Therefore, based on Eq. (11), $M_{\boldsymbol{\kappa}}$ retains six unit eigenvalues (i.e., periodicity of $\boldsymbol{\kappa}(t)$) for an arbitrary linear invertible map $\Psi(t)$. Thus, a set of invariant integration constants $\boldsymbol{\kappa}(t)=\boldsymbol{\kappa}$ can be defined with $\Psi(t)$ being a fundamental matrix solution to the variational equation, which leads to the well-known ROE theory [5] based on two periodic orbits, as shown in Fig. 3(a). Note that this eigensystem provides five degrees of freedom in the design space of bounded relative orbits.
@@
When additional perturbations are considered (i.e., PR2BP), some eigenvalues of the monodromy matrix perturb from unity. Due to the additional conditions required for periodicity, the solution set for periodic relative motion is reduced. Previous works address bounded relative orbit designs that cancel a secular drift caused by the $J_{2}$-perturbation [5, 54].
@@ 3.1.2 Eigensystem in the RMBP
The local linearization of the dynamics of RMBPs contains time-varying components [19], preventing a closed-form expression of the monodromy matrix. Instead, the monodromy matrix can be numerically integrated (cf. Eq. (9)) concurrently with the periodic solution (cf. Eqs. (39a), (41a), and (42a)) for the time interval $t_{0},t_{0}+T]$, using the plant matrix $\boldsymbol{A}(t)$. This system becomes a 42-dimensional ordinary differential equation.
Generally, not all eigenvalues of the monodromy matrix computed along a periodic solution of RMBP are unity. Therefore, the entire solution set of bounded relative motion around a periodic orbit within the RMBPs is severely constrained. One trivial solution is a set of relative motions between spacecraft on the same periodic orbit but with phase offsets. Besides, a periodic orbit in the RMBPs may admit up to two oscillatory modes in the corresponding monodromy matrix, characterized by a conjugate pair of complex unimodular eigenvalues [17, 40]. When the oscillatory mode is excited, an initial relative state within the corresponding eigenspace evolves into a quasi-periodic motion; after one period $T$, the relative position and velocity do not repeat but remain confined within a rotating ellipse, known as an invariant circle. This structure is illustrated in Fig.  2(b) as dotted ellipses on the eigenspace.
@@
Figure 3: Periodic and quasi-periodic motion in the R2BP and RMBP.
A QPRO constructed based on the oscillatory mode in the linearized dynamics is first-order equivalent to a QPO in the synodic frame that envelopes a reference orbit under the original nonlinear dynamics. The QPO evolves along the surface of a two-dimensional Quasi-Periodic Invariant Torus (QPIT), which constitutes a nonlinearly stable geometric structure in the phase space of the RMBPs. This torus exhibits bounded perturbation responses in transverse directions and preserves its quasi-periodic motion under small disturbances. This enables solutions to evolve along it without altering their overall shape, as shown in Fig. 3(b). This bounded motion is ensured because each point on the torus surface follows the vector field of the system’s equations of motion. Any position and velocity ($\mathbb{R}^{6}$) on the surface of a QPIT can be characterized by a two-dimensional phase vector $\boldsymbol{\theta}\in\mathbb{T}^{2}$. Consequently, a motion initiated from a point on the torus can be viewed as an evolution of $\boldsymbol{\theta}$. The progression along a periodic orbit of the RMBPs is characterized by $\theta_{0}$, whereas $\theta_{1}$ represents the phase within an invariant curve, defined as a closed curve where quasi-periodic solutions are confined under the $T$-stroboscopic map. In this paper, a QPIT resolved in the local frame around the chief is termed Quasi-Periodic Relative Invariant Torus (QPRIT). Compared to the QPIT around the attractor, the structure of the QPRIT around the chief spacecraft provides a fundamental geometry of a QPRO, enabling a simple characterization of the relative motion as shown in the following sections. Thus, an invariant circle is equivalent to the first-order approximation of the invariant curve (cf. Fig. 3(b)) in a local frame [28], leading to two degrees of freedom in the QPRO design in each oscillatory mode.
@@ 3.2 Local Toroidal Coordinates
Local Toroidal Coordinates (LTC) [40] are coordinate systems on a local frame along a reference orbit that project the dynamics onto the oscillatory eigenspace of the RMBPs. Instead of a full linear mapping to integration constants, a mode-isolating transformation is employed to emphasize the dynamical mode that provides a quasi-periodic bounded relative motion. In this paper, it is assumed that the chief is on a periodic orbit of RMBPs that contains at least one oscillatory mode.
@@
$\boldsymbol{\psi}(t)=\varepsilon\left(\boldsymbol{w}_{r}(t)\cos\theta+\boldsymbol{w}_{i}(t)\sin\theta\right),\quad\boldsymbol{w}_{r}=\begin{bmatrix}\boldsymbol{r}_{r}(t)\\ [\dot{\boldsymbol{r}}_{r}(t)]_{\mathcal{V}}\end{bmatrix},\boldsymbol{w}_{i}=\begin{bmatrix}\boldsymbol{r}_{i}(t)\\ [\dot{\boldsymbol{r}}_{i}(t)]_{\mathcal{V}}\end{bmatrix}$, (12)
Figure 4: Nonsingular local toroidal coordinates $\alpha,\beta,h]$ and geometric local toroidal coordinates $\varepsilon,\theta,h]$
@@
With this in mind, the local toroidal frame is defined based on a time-varying non-orthogonal and non-unimodular basis triad $\{\boldsymbol{r}_{r}(t),\boldsymbol{r}_{i}(t),\hat{\boldsymbol{n}}(t)\}$, where $\hat{\boldsymbol{n}}(t)$ is a unimodular vector parallel to $\boldsymbol{r}_{r}(t)\times\boldsymbol{r}_{i}(t)$. In order to achieve a unique and consistent representation of an eigenvector and to construct a triad from a local eigensystem, basis vectors at the fixed point at the initial epoch $\{\boldsymbol{r}_{r}(t_{0}),\boldsymbol{r}_{i}(t_{0})\}$ are normalized and rotated via singular value decomposition (SVD) so that these vectors are orthogonal and $\boldsymbol{r}_{r}(t_{0})$ becomes unimodular. For details of the normalization process, see Ref. 39. Once normalized at the fixed point, the evolution of $\boldsymbol{w}$ can be propagated by the linear dynamics
$[\dot{\boldsymbol{w}}(t)]_{\mathcal{V}}=\boldsymbol{A}(t)\boldsymbol{w}(t),$ (13)
where the plant matrix is obtained from Eq. (8), and the time derivative is taken in the local frame. Therefore, the basis vectors $\{\boldsymbol{r}_{r}(t),\boldsymbol{r}_{i}(t),\hat{\boldsymbol{n}}(t)\}$ are generally neither orthogonal (nor unimodular for $\{\boldsymbol{r}_{r},\boldsymbol{r}_{i}\}$) except for the initial triad at the fixed point.
Within the local toroidal frame, the relative position of the deputy to the chief is written as $\boldsymbol{\rho}=\alpha\boldsymbol{r}_{r}+\beta\boldsymbol{r}_{i}+h\hat{\boldsymbol{n}}$. Based on this, nonsingular LTC is defined as $\boldsymbol{\zeta}_{\text{ns}}:=[\alpha,\beta,h,\dot{\alpha},\dot{\beta},\dot{h}]^{\top}$, where time derivatives are taken in the local (toroidal) frame. Using Eq. (10), the mapping from the nonsingular LTC $\boldsymbol{\zeta}_{\text{ns}}$ to the relative position and velocity resolved in the local frame $\boldsymbol{x}$ is related as
$\boldsymbol{\zeta}_{\text{ns}}=T(t)^{-1}\boldsymbol{x},\quad T(t):=\begin{bmatrix}R(t)&\boldsymbol{0}_{3}\\ R^{\prime}(t)&R(t)\end{bmatrix}\Leftrightarrow T^{-1}(t)=\begin{bmatrix}R(t)^{-1}&\boldsymbol{0}_{3}\\ -R(t)^{-1}R^{\prime}(t)R(t)^{-1}&R(t)^{-1}\end{bmatrix}$, (14)
where $R(t):=[\boldsymbol{r}_{r}(t),\boldsymbol{r}_{i}(t),\hat{\boldsymbol{n}}(t)]$ and $R^{\prime}(t):=[[\dot{\boldsymbol{r}}_{r}(t)]_{\mathcal{V}},[\dot{\boldsymbol{r}}_{i}(t)]_{\mathcal{V}},[\dot{\hat{\boldsymbol{n}}}(t)]_{\mathcal{V}}]$. The time derivative of the unit normal vector $\hat{\boldsymbol{n}}(t)$ is computed as
$[\dot{\hat{\boldsymbol{n}}}(t)]_{\mathcal{V}}=[\dot{\boldsymbol{r}_{r}}(t)]_{\mathcal{V}}\times\boldsymbol{r}_{i}(t)+\boldsymbol{r}_{r}(t)\times[\dot{\boldsymbol{r}_{i}}(t)]_{\mathcal{V}}-(\boldsymbol{n}(t)^{\top}([\dot{\boldsymbol{r}_{r}}(t)]_{\mathcal{V}}\times\boldsymbol{r}_{i}(t)+\boldsymbol{r}_{r}(t)\times[\dot{\boldsymbol{r}_{i}}(t)]_{\mathcal{V}}))\boldsymbol{n}(t).$ (15)